mathslover (mathslover):

Question : A ball of mass 0.3 kg coming towards a player with a velocity of 10 ms–1 is kicked back with a velocity of 20 ms–1 in the opposite direction. If the impact lasts for 1/30 th of a second, what is the average force involved?

mathslover (mathslover):

I got 90 N somehow :P but the answer is wrong

mathslover (mathslover):

I took u = 10 and v = 20

OpenStudy (xishem):

Did you use the impulse equation? \[\vec I=\sum \vec F \Delta t\]

mathslover (mathslover):

and t = 1/30 so what we have : a = (v-u)/t = (10)/(1/30) = 300 m/s^2

mathslover (mathslover):

F = 0.3 kg * 300 m/s^2 = 90 N

mathslover (mathslover):

@Xishem that is absolutely NOT in our course.. so this will not be used here... any easier method like I did?

OpenStudy (xishem):

Hmmm. One sec.

mathslover (mathslover):

oh k

OpenStudy (shane_b):

Why not use the momentum-impulse theorem?\[mv_f-mv_f=Ft\]

OpenStudy (xishem):

Your problem is that you aren't accounting for the fact that the second velocity is in the negative x direction.

OpenStudy (shane_b):

Typo fixed:\[mv_f-mv_i=Ft\]

OpenStudy (xishem):

You can use F=ma where a is the average acceleration. You'll get:\[F=ma \rightarrow F=m(\frac{v_f-v_i}{\Delta t})\]

mathslover (mathslover):

right

OpenStudy (xishem):

That's what you did, just make sure you use the correct signs, and you end up getting:\[F=(0.3kg)(\frac{(-20\frac{m}{s}-10\frac{m}{s})}{\frac{1}{30}s})=-270N\]

OpenStudy (shane_b):

Yep ^^

mathslover (mathslover):

oh why -ve there?

OpenStudy (xishem):

I don't think he wants to use impulse theory because I don't think they've covered momentum yet?

OpenStudy (shane_b):

@mathslover: Because the ball changed directions...

OpenStudy (shane_b):

@Xishem: You and I both used the same formula :)

OpenStudy (xishem):

The initial velocity is 10m/s in +x direction. It hits a wall and reverses direction to the -x direction at 20m/s. The total change in velocity is -30m/s.

OpenStudy (xishem):

Yeah, it breaks down into the same formula, but mine doesn't require an understanding of momentum or impulse :P.

OpenStudy (xishem):

Isn't the entirety of physics like that?

mathslover (mathslover):

right thanks a lot @Xishem and @Shane_B @Xishem can u please give a medal to shane_B since m not able to give medal to both

OpenStudy (shane_b):

My crystal ball says his next chapter will cover impulse and momentum :)

mathslover (mathslover):

Well I had studied momentum : p = m v and the : momentum before collision = momentum after collision like this ... but just forgot about "impulse" I got this now thanks a lot again

OpenStudy (xishem):

Perhaps. Maybe it will be circular motion, though :P. Or energy. Or... etc.

OpenStudy (shane_b):

:)

OpenStudy (xishem):

Or maybe... relativity :D

mathslover (mathslover):

the answer is "270 N " why not -270?

OpenStudy (xishem):

It depends on your frame of reference. Either answer should be acceptable. Make sure you draw a diagram of the situation which indicates the direction of the x-axis to clarify. As long as the force is directed in the opposite direction of the initial velocity, you're okay.

mathslover (mathslover):

oh k got it thanks again :)

OpenStudy (shane_b):

The bottom line is the magnitude of the force = |F|=|-270N|=270N

mathslover (mathslover):

right ~~